1000 ml buffer is 0.175 M of HClO and 0.150M in HClO and the pH is 7.47

1. what is the pH when 10.00ml of 2.0M NaOH is added?

2. if you add 150mg of HBr what will be the new pH (assume no volume change). Molar mass of HBr is 80.91g/mol

sorry HClO it is 0.175 M and for NaClO it is 0.150
Answer #1

[base] PH - pКа +10g [acid] [CIO 7.47 pKa+log [HClo] 0.150 7.47 pKa+log - 0.175 pKa -0.06695 pKa 7.470.06695 = 7.537 1

(a) Number of moles of HCIO = M*V = 0.175 M* 1000 mL 175 mmol Number of moles of NaClO M*V 0.150 M * 1000 mL 150 mmol Number

(b) Number of moles of HCIO = M*V = 0.175 M* 1000 mL 175 mmol Number of moles of NaClO M*V = 0.150 M * 1000 mL 150 mmol Numbe